Kofinatland is a hypothetical economy with the following economic indicators:
-Size of population:8million
-GDP at market price(-GDP at m.pm.)
-Subsidies: US $50 billion
-Interest paid by consumers:US $27billion
-Employees contribution to social insurance: US $200 billion
-Net factor income from abroad(NFYFA) :US $200 billion
-Capital Consumption allowance(CCA):US $80billion
-Human development index(HDI):0.91
-Indirect taxes on goods and services(ITGS):US $120billion
-Life expectancy rate:87years
-Government transfer payment: US $71.5 billion
-Factor income of foreigners in the economy(FYF): US $80 billion
-National Population growth rate: 0.5%

Determine the economy's
i. Net national product at market cost (NNP at m.px)
ii. Net national product at factor cost (NNP at f.c)
iii. Factor income of national abroad (FYN)
iv.GDP per capita ( note to use the -GDP at m.px. in your calculation)
1
Expert's answer
2020-05-12T10:48:09-0400

i.Net National Product at Market Prices (NNPMP) = GNPMP – Depreciation=GDP at m.pm.-Net factor income from abroad-Capital Consumption allowance=570-200-80=290

ii. Net national product at factor cost=GDP at m.pm.-net indirect taxes-Capital Consumption allowance=570-120-80=370

iii. .Factor income of nationals abroad (FYN)=NDP at MP +Capital Consumption allowance + Factor income from abroad - GNP at FC - Indirect Taxes + Subsidies=570-80+80+200-570-120+50=130

iiii."\frac{570}{8}=71.25"